# Sparsity-Aware Coding for Single Photon Sensitive Vision Using Selective Sensing

### v4

@@ Abstract
Optical coding is widely used in computational imaging systems and is a good approach for designing vision systems. However, most coding methods are developed assuming additive Gaussian noise, while modern optical imaging systems are mainly affected by Poisson noise. Previous studies have highlighted the significant differences between these noise models and proposed coding optimization algorithms for image recovery under Poisson noise. They concluded that the compressibility arising from data variance is crucial for image recovery under Poisson noise. This makes a strong case for the design of end-to-end vision systems that avoid image formation, since the data-driven vision tasks, typically downstream of imaging, is more compressible than imaging itself. In this project, we propose a coding strategy by jointly optimizing an entire vision system, including measurement and inference, using the classification accuracy as a metric. We demonstrate the importance of incorporating Poisson noise in optimizing even the simplest vision systems and propose an approach to achieve it.
@@
Optical coding, Poisson noise, single-pixel-camera, data-driven prior, optical-neural-networks
@@ I Introduction
Optical coding, also known as multiplexing, is a popular technology in computational imaging (CI) due to its potential to achieve enhanced signal-to-noise ratio (SNR) and greater light throughput than point by point measurements[1]. If ignoring the noise, coding is represented as
@@
where $\boldsymbol{x}$ is the signal vector, $\boldsymbol{M}$ is the sensing matrix, and $\boldsymbol{y}$ is the measured vector. Decoding refers to the process of recovering the initial signal $\boldsymbol{x}$ from $\boldsymbol{y}$. The recovery quality of $\boldsymbol{x}$, however, also depends on the conditioning of the $\boldsymbol{M}$ besides the light throughput [1]. Suboptimal $\boldsymbol{M}$ have the potential to significantly degrade the performance of the vision system during the recovery process, as highlighted by Mitra et al. [1]. The optimization of coding schemes regarding $\boldsymbol{M}$ is intricately linked to the characteristics of noise. Coding schemes, typically formulated under the assumption of signal-independent Additive Gaussian Noise (AGN), offer a mathematically tractable framework. However, AGN, while suitable for modeling certain noise types introduced by measurement devices, inadequately represents Poisson noise [2] arising intrinsically due to the discrete and random nature of photon measurements. Past works usually distinguish between low-light regimes, where AGN introduced by imperfect sensors dominates, and bright-light regimes, where the physical photon noise dominates [1]. With advances in sensor design, even the sensor noise of low-end camera sensors has been reduced to a few photons per pixel. This reduction is significant enough that photon noise becomes the primary noise factor in the vast majority of imaging applications. It has been shown that in that case without the use of priors for regularization, coding does not provide an improvement over a simple sequential scan [3, 4, 5, 6, 7]. Mitra et. al. [1] develop a data driven prior to design optimized codes for image reconstructions that can markedly enhance the reconstruction performance of coding schemes under Poisson noise [1]. An important conclusion is that if the signal exhibits sparsity and thereby compressibility, signal optimized coding can improve recovery performance under Poisson noise and the amount of improvement is directly related to the comprehensibility of the measured data. However, recovery is typically not the ultimate goal for modern vision systems. Since vision tasks, such as classification, operate in embeddings created from images their data should be inherently more compressible than their upstream imaging tasks. Therefore, if signal compressibility is the key, skipping the intermediate imaging step and directly focusing on coding for better feature extraction for subsequent vision tasks should yield better performance. To show this, our study implements a simple end to end vision system that can be optimized under different noise assumptions. We illustrate why AGN is not suitable for optimizing coding under Poisson noise and show that a coding scheme optimized for a specific vision task and for the correct noise model can perform close to the optimal non-coding camera even when only photon noise is present.
@@
  • We introduce the methodology of Selective Sensing (SS), which encompasses coding techniques specifically designed to extract data-driven priors selectively. These coding methods can be seamlessly integrated with models for downstream tasks, providing a novel approach to efficiently incorporate learned priors into various applications.
@@ II Related Work
Coding under Poisson noise. In coded imaging, Hadamard matrices are considered the optimal coding scheme for multiplexing [3, 8, 9, 1] in systems with only AGN, though this conclusion doesn’t hold for Poisson noise [1]. In many recent projects, optimization efforts focus on algorithmic enhancements for reconstruction, often maintaining the use of random coding strategies [10, 11, 12, 13], though this type of matrices are not recommended under Poisson noise[5, 14, 7]. Wuttig, Ratner, et al. [9, 15] introduced Poisson noise-aware optimization techniques for multiplexing matrices, outperforming Hadamard in the presence of moderate signal-dependent noise. However, their analysis neglects data priors, leading to suboptimal matrices for greater Poisson noise [1]. Additionally, attempts at matrix optimization focusing on minimal mutual coherence [16] lack a foundation in the Poisson noise model, as highlighted in this study.
@@
Feature Specific Imaging. Feature-specific imaging (FSI) is a type of imaging system where sensing matrix $\boldsymbol{M}$ incorporates linear features of expected data [17, 18]. This approach can provide higher feature fidelity and lower detector count than conventional imaging, especially for applications that require relatively few features [17, 18, 19]. This technique can be viewed as a variant of compressed sensing, wherein the sensing matrix is determined based on prior information [19]. Nevertheless, current state-of-the-art FSI is explicitly designed for AGN rather than Poisson noise[19]. FSI can be regarded as a technique that leverages data driven priors, and Mitra et al. have asserted that employing data driven priors can enhance image recovery under Poisson noise [1]. Therefore, in our work, we underscore the promising prospects of extending feature-specific concepts to a broad spectrum of vision tasks under Poisson noise.
@@
End-to-end optimization. This method jointly optimizes camera design and digital signal processing to achieve optimal overall performance [20, 21, 22]. In many previous projects, this idea was usually implemented without considering Poisson noise in cost functions [23, 24, 25, 26, 27, 28, 22, 11] or without optimizing masks with gradient incorporating Poisson distribution [29, 20, 30, 31, 32, 11, 33]. Rego et. al froze the sensing matrix as a pinhole without optimizing it [30]. Wang et al. [32] successfully implemented a neural network model for handwritten number classification on an optical device with limited photon budget, demonstrating the potential for AI-assisted optimization of coding schemes. However, the Poisson noise was considered only in model testing where the most robust model was picked among a set of different initialization seeds [32]. Xu et al. [33] examined the robustness of the proposed neural networks model to Poisson noise; however, the training process does not explicitly incorporate the gradients affected by Poisson noise. Our contribution is developing a noise-aware training approach within a neural network model to find the globally optimal masks under Poisson noise. Addressing the limitations identified by Mitra et al. [1], this model presents the following advantages.
  • The optimization of coding finds a globally optimal $\boldsymbol{M}$ rather than its greedy approximation.
  • Noise is incorporated during the training process to simulate real-world conditions instead of using a one-dimensional affine noise model.
  • The training process is accelerated and computationally more efficient due to the binary constraint being addressed even in real experiments. Additionally, the prior is learned from the entirety of the light fields and spectral data, rather than from patches.
  • The contribution of a data-driven prior to performance enhancement is more accurately discerned using the classification rate as the metric.
@@ III-A Single Pixel Camera
The single-pixel camera shown in Fig. 1 is a popular example of coding, and serves as the primary example in this project. It employs a rapid single-pixel detector in conjunction with a Digital Micromirror Device (DMD) to sequentially multiplex a sensing matrix by modulating various patterns on the DMD [1]. Its measurement process mathematically follows the Eq. 1. Here, $\boldsymbol{x}\in\mathbb{R}^{N\times 1}$ is the image representation of the field of view (FOV), which consists of $N$ pixels, and is measured by linear projections, or masks, $\boldsymbol{M}\in\mathbb{R}^{m\times N}$ to attain corresponding flux levels or photon counts, $\boldsymbol{y}\in\mathbb{R}^{m\times 1}$ [5]. Physically, $\boldsymbol{M}$ can be implemented on the DMD by directing or blocking different parts of the incoming light from $\boldsymbol{x}$ [34] and averaging them on sensors that digitize the detected flux levels or photon counts $\boldsymbol{y}$. If the sensing matrix $\boldsymbol{M}$ is full-rank, the $\boldsymbol{x}$ can be reconstructed by $\boldsymbol{M}^{-1}\boldsymbol{y}$. However, even when $m<N$, we can still recover the $\boldsymbol{x}$ via compressed sensing. While Hadamard matrices are considered effective for $\boldsymbol{M}$, their suitability diminishes in the presence of data-dependent noise [1, 3]. In this work, we will study the performance of this camera for different choices of $\boldsymbol{M}$ under different noise models.
@@ III-B Selection of Sensing Matrix
Eq. 1 is the foundational equation for single-pixel camera. Based on it, we distinguish different types of $\boldsymbol{M}$ below for reference and guidance.
@@ III-B1 Coding Types
  • Impulse imaging entails an ideal camera that simultaneously captures all elements in the measurements, $\boldsymbol{y}$. A prominent instance is a pixel array. When the field of view $\boldsymbol{x}$ consists of $N$ pixels, the sensing matrix assumes the form $\boldsymbol{M}=N\boldsymbol{I}$ where $\boldsymbol{I}$ is an identity matrix. Each pixel receives $N$ times the exposure in comparison to the raster scan scenario. This technique sets a benchmark akin to a gold standard, serving as an upper limit for the CI performance and an indicator of the noise level for comparison.
@@ III-B2 Vision Tasks
  • Computational Imaging (Weak-prior/Some Compression): In this kind of tasks, a weak-prior assumption is made. This idea, which focuses on image reconstruction, is based on prior knowledge, such as signals being sparse in some linear space, with noise concentrated in the high-frequency domain. One example in this category is compressed sensing.
  • Vision (Strong-prior/Much Compression): Many computer vision tasks involve a strong prior, enabling substantial compression during measurements. Directly capture of principal features by $\boldsymbol{M}$ is possible if some of its components resemble the $\boldsymbol{x}$. In these tasks, a few measurements are often sufficient with a strong prior derived from training data. Our focus is to optimize codes in this scenario.
@@ III-C Constraints of sensing matrices
The implementation of sensing matrix $\boldsymbol{M}$ usually involves the following constraining factors.
@@
    Positivity-preserving [5]. It is not possible to physically implement negative values for the masks $\boldsymbol{M}$, as demonstrated in [18, 5]. In this project, we employed the dual-rail approach as outlined in the work of Neifeld et al where the positive and negative entries of a mask give rise to two separate measurements [18].
@@ III-D Model under Noise
This project investigates two noise models. The first one is the AGN related to dark current or read noise from imperfect sensor materials, which mainly originates from thermal vibrations of atoms at sensors. In this noise model, Eq. 1 becomes:
@@
where $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\sigma^{2}\boldsymbol{I})$, and $\sigma$ is the standard deviation. The other noise model is the Poisson noise or shot noise, which arises from the statistical nature of photons [35]. The measurement equation for this noise model is:
@@
where $y_{i}=\sum_{j=1}^{N}M_{ij}x_{j}$ [5]. It should be noted that the constraint 2 in Eq. 3 prohibits negative entries in $\boldsymbol{M}$.
@@ IV-A Model Configuration
All coding strategies in this project for classification tasks share the same fundamental model structure, as illustrated in Fig. 2. In this project, we extend the applicability of our model to handwritten number classification with the MNIST dataset and classification of hyperspectral data. To classify handwritten numbers using optical devices with limited photon budget, we built a PyTorch-based pipeline consisting of a scanner module including $\boldsymbol{M}$ and a noise generator to simulate the acquisition of $\tilde{\boldsymbol{y}}$ described in Eq. 2 and Eq. 3, and a classifier module classifying the measured objects by $\tilde{\boldsymbol{y}}$. For the classification task, we utilized a two-hidden layer artificial neural network (ANN) with 40 and 128 nodes, serving as the primary model. To prevent overfitting and enhance generalizability, we applied dropout after each hidden layer, followed by a rectified linear unit (ReLU) activation layer. Once operational, this model can serve various purposes, such as object classification and signal recovery. For instance, it can process images of handwritten numbers, providing the likelihood of each image corresponding to a specific number by cross-entropy loss. Furthermore, the classifier can function as a decoder to recover the input image by MSE loss, which is adopted in section IV-B.
During the training of the model, the scanner computes the photon counts with noise and then transmits this data to the classifier. The noise is sampled from a pre-defined noise distribution and varies across epochs. Once the loss has been computed, gradient descent is used to optimize the parameters of the classifier. In addition, we propose an end-to-end optimization model called Optical Neural Networks (ONN). What sets the ONN apart from other coding schemes is the learnable $\boldsymbol{M}$ which is not predefined or fixed. In other words, back-propagation is extended into the scanner, allowing for the simultaneous optimization of the masks and the classifier under different noise models. This approach offers distinct advantages compared to traditional methods, which typically optimize the classifier independently of the scanner or optimize the scanner without the appropriate noise model.
@@
Optimizing the scanner can be challenging as the Poisson noise generating function typically lacks a gradient. While reparameterization is a useful method for addressing AGN, it is not as effective for addressing Poisson noise. To address this limitation, we employ a method mentioned by Cossairt et al. that utilizes a normal distribution with a variance equal to its expectation to approximate the Poisson distribution during training [8]. This approximation enables us to employ PyTorch builtin reparameterization method, i.e. rsample, and optimize the scanner more effectively under Poisson noise.
To approximate the Poisson noise by AGN, we can re-write the measurement with noise $\tilde{\boldsymbol{y}}=\boldsymbol{Mx}+\boldsymbol{J}$ where $\boldsymbol{J}=(\boldsymbol{Mx})^{\circ\frac{1}{2}}\odot\boldsymbol{\epsilon},\boldsymbol{\epsilon}\sim\mathcal{N}(0,\boldsymbol{I})$ [36]. During the training under Poisson noise, $\boldsymbol{J}$ also plays a role in gradient computation, which is the most significant difference compared with training under AGN.

### v8

# An approach to Physics Based Single Photon Vision System Design

@@ Abstract
The design of the camera and optical measurement is a crucial part of optimizing machine vision systems. However, camera designs are usually optimized to produce human-interpretable images. Moreover, camera optimization typically makes the assumption of additive noise, while modern optical imaging systems are mainly affected by photon noise which is not additive. Previous studies have highlighted the fundamental effect of the noise model on the outcome of the design process in the context of coded single-pixel and compressed sensing cameras.
In addition to the exact noise model, the nature of the data encountered and the vision task fundamentally affect the optimal hardware design. This means that a camera or compressed sensing code that is optimal for capturing images is almost certainly not optimal for any downstream vision task. In this work, we used a simple end-to-end model on a single-pixel camera to study the effects of the noise model for different vision tasks. We demonstrate the importance of incorporating non-linear noise features in optimizing even the simplest vision systems.
@@
Optical coding, photon noise, single-pixel-camera, data-driven prior, optical-neural-networks
@@ I Introduction
While we can mass-produce optical components of increasing complexity, most vision hardware still closely mimics the optics of the human eye to create easily interpretable images. In this approach, the lens projects an image that is then post-processed to extract patterns. However, more generalized vision systems can perform alternative measurements that may be more perceptually meaningful than simply producing images. Designing such systems requires a collaborative effort to integrate hardware and digital processing algorithms while considering the appropriate noise model. The significance of this approach is particularly illustrated by the single-pixel camera in this context.
The single pixel camera uses the concept of optical coding, or multiplexing, which involves projecting an image onto a mask and collecting the transmitted light with a large sensor pixel. The system can potentially achieve an improved signal-to-noise ratio (SNR) and higher light throughput by coding compared to point-by-point measurements [1]. In this context, a measurement vector $\boldsymbol{y}$ is formed from a set of measured photon counts, with the corresponding masks vectorized as rows of the sensing matrix $\boldsymbol{M}$. Ignoring the noise, coding can be represented as
@@
where $\boldsymbol{x}$ is the signal vector. Decoding refers to the process of recovering the initial signal $\boldsymbol{x}$ from $\boldsymbol{y}$. The recovery quality of $\boldsymbol{x}$ is significantly dependent on the conditioning of $\boldsymbol{M}$ and the light throughput of $\boldsymbol{M}$ [1]. Suboptimal $\boldsymbol{M}$ have the potential to significantly degrade the performance of the vision system during the recovery process, as highlighted by Mitra et al. [1].
While a greater light throughput, or a lower noise level, can improve recovery, the improper choice of the noise model can also give rise to a suboptimal coding at specific noise levels. Coding design is often performed under the assumption of signal-independent additive Gaussian Noise (AGN) which arises from imperfect sensors. With advances in sensor design, even the sensor noise of low-end camera sensors has been reduced to a few photons per pixel. This reduction is significant enough so that nonadditive photon noise becomes the primary noise factor in the vast majority of imaging applications [2]. Photon noise follows a super-Poissonian distribution [3]. For passive cameras operated with incoherent ambient light the noise can usually be approximated as Poissonian.
The transition from AGN to more realistic photon noise fundamentally alters the designs of vision systems. For example, although single-pixel cameras with random masks provide performance advantages under AGN, they do not outperform a simple sequential scan when dealing with Poisson noise [4, 5, 6, 7, 8]. Moreover, optimal coding designs under AGN conditions do not deliver optimal performance under Poisson noise at the same noise levels.
To address this issue, Mitra et al. [1] develop a method using data-driven priors to design optimized codes for image reconstructions. Their approach significantly improves the performance of coding under Poisson noise by designing measurements specifically targeting sparsity or compressibility in the data [1]. However, image recovery is typically not the ultimate goal for modern vision systems. Since vision tasks, such as classification, operate in embeddings created from images, their data should be inherently more compressible than their upstream imaging tasks, allowing for more effective measurement designs. To show this, we implement a simple end-to-end vision system that can be optimized under different noise assumptions. We illustrate why AGN is not suitable for optimizing coding under photon noise and show that a coding scheme optimized for a specific vision task and for the correct noise model can perform close to the optimal non-coding camera even when photon noise is present.
@@
  • We introduce the methodology of Selective Sensing (SS), which encompasses coding techniques specifically designed to extract data-driven priors selectively for various downstream tasks.
@@
  • We propose a hybrid optimization method for non-differentiable Poisson noise that employs AGN with signal-dependent variance and the reparameterization trick to estimate gradients during training, while switching to Poisson noise in the testing stage for model assessment.
Our model is extendable to more general camera optics and different vision tasks in future work.
@@ II Related Work
Coding under photon noise. In coded imaging, Hadamard matrices are considered the optimal coding scheme for multiplexing [4, 2, 9, 1] in systems with only AGN, although this conclusion does not hold for photon noise [1]. In many recent projects, optimization efforts focus on algorithmic enhancements for reconstruction, often maintaining the use of random coding strategies [10, 11, 12, 13], though this type of matrices are not recommended under photon noise[6, 14, 8]. Wuttig and Ratner, et al. [9, 15] introduced photon-noise-aware optimization techniques for coding matrices that outperform Hadamard in the presence of moderate signal-dependent noise. However, their analysis neglects the data priors, leading to suboptimal matrices for greater photon noise [1]. Additionally, attempts at matrix optimization focusing on minimal mutual coherence [16] do not adequately consider photon noise, as highlighted in this study.
@@
Feature Specific Imaging. Feature-specific imaging (FSI) is a type of imaging system where sensing matrix $\boldsymbol{M}$ incorporates linear features of expected data [17, 18]. This approach can provide higher feature fidelity and lower detector count than conventional imaging, especially for applications that require relatively few features [17, 18, 19]. This technique can be viewed as a variant of compressed sensing, in which the sensing matrix is determined based on prior information [19]. However, current state-of-the-art FSI is explicitly designed for AGN rather than Poisson noise[19]. FSI can be regarded as a technique that leverages data driven priors, and Mitra et al. have asserted that employing data driven priors can enhance image recovery under Poisson noise [1]. Therefore, in our work, we underscore the promising prospects of extending feature-specific concepts to a broad spectrum of vision tasks under Poisson noise.
@@
End-to-end optimization. This method jointly optimizes camera design and digital signal processing to achieve optimal overall performance [20, 21, 22]. In many previous projects, this idea was usually implemented without considering Poisson noise in cost functions [23, 24, 25, 26, 27, 28, 22, 11] or without optimizing masks with gradient incorporating Poisson distribution [29, 20, 30, 31, 32, 11, 33]. Rego et. al froze the sensing matrix as a pinhole setting without optimizing it [30]. Wang et al. [32] successfully implemented a neural network model for handwritten number classification on an optical device with limited photon budget, demonstrating the potential for AI-assisted optimization of coding schemes. However, Poisson noise was considered only in model testing where the most robust model was picked among a set of different initialization seeds [32]. Xu et al. [33] examined the robustness of the proposed neural network model to Poisson noise; however, the training process does not explicitly incorporate the gradients affected by Poisson noise. As we see later, however, the correct noise model is essential to obtain an optimal or near-optimal optical design. Our contribution in this area is developing a noise-aware training approach within a neural network model to find globally optimal masks under Poisson noise. Although our approach also employs data-driven priors, which is similar to that of Mitra et al., our research extends beyond simple image reconstruction to address a wider spectrum of vision tasks and investigates task-specific optimization techniques on optical hardware.
@@ III-A Single Pixel Camera
The single-pixel camera shown in Fig. 1 is a popular example of coding, and serves as the primary example in this project. It employs a rapid single-pixel detector in conjunction with a Digital Micromirror Device (DMD) to sequentially multiplex a sensing matrix by modulating various patterns on the DMD [1]. Its measurement process mathematically follows the Eq. 1. Here, $\boldsymbol{x}\in\mathbb{R}^{N\times 1}$ is the image representation of the field of view (FOV), which consists of $N$ pixels, and is measured by linear projections or masks, $\boldsymbol{M}\in\mathbb{R}^{m\times N}$ to attain corresponding flux levels or photon counts, $\boldsymbol{y}\in\mathbb{R}^{m\times 1}$ [6]. Physically, $\boldsymbol{M}$ can be implemented on the DMD by directing or blocking different parts of the incoming light from $\boldsymbol{x}$ [34] and averaging them on sensors that digitize the detected flux levels or photon counts $\boldsymbol{y}$. If the sensing matrix $\boldsymbol{M}$ is full-rank, the $\boldsymbol{x}$ can be reconstructed by $\boldsymbol{M}^{-1}\boldsymbol{y}$. However, even when $m<N$, we can still recover the $\boldsymbol{x}$ via compressed sensing. While Hadamard matrices are considered effective for $\boldsymbol{M}$, their suitability diminishes in the presence of data-dependent noise [1, 4]. In this work, we will study the performance of this camera for different choices of $\boldsymbol{M}$ under different noise models.
@@ III-B Selection of Sensing Matrix
Eq. 1 is the foundational equation for a single-pixel camera. Based on it, we distinguish different types of $\boldsymbol{M}$ below for reference and guidance.
@@ III-B1 Coding Types
  • Impulse imaging (conventional imaging) entails an ideal camera that simultaneously captures all elements in the measurements, $\boldsymbol{y}$. The key feature of this system is its use of a pixel array sensor instead of a single-pixel sensor. This allows for the detection of all incoming photons, whereas the raster scan only collects a small fraction, specifically $\frac{1}{N}$ when the field of view $\boldsymbol{x}$ consists of $N$ pixels. Therefore, the sensing matrix assumes the form $\boldsymbol{M}=N\boldsymbol{I}$ where $\boldsymbol{I}$ is an identity matrix. This technique sets a benchmark akin to a gold standard, serving as an upper limit for the CI performance and an indicator of the noise level for comparison.
@@ III-B2 Vision Tasks
  • Computational Imaging (Weak-prior/Some Compression): In this kind of task, a weak-prior assumption is made. This idea, which focuses on image reconstruction, is based on prior knowledge, such as signals being sparse in some linear space, with noise concentrated in the high-frequency domain. One example in this category is compressed sensing.
  • Vision (Strong-prior/Much Compression): Many computer vision tasks involve a strong prior, enabling substantial compression during measurements. Activation of the main feature directly by $\boldsymbol{M}$ is possible if some of its components resemble $\boldsymbol{x}$. In these tasks, a few measurements are often sufficient with a strong prior derived from training data. Our focus is on optimizing the codes in this scenario.
@@ III-C Constraints of sensing matrices
The implementation of the sensing matrix $\boldsymbol{M}$ usually involves the following constraining factors.
@@
    Positivity-preserving [6]. It is not possible to physically implement negative values for masks $\boldsymbol{M}$, as demonstrated in [18, 6]. In this project, we used the dual-rail approach as outlined in the work of Neifeld et al. where the positive and negative entries of a mask give rise to two separate measurements [18].
@@ III-D Model under Noise
This project investigates two noise models. The first one is AGN related to dark current or read noise from imperfect sensor materials, which mainly originates from thermal vibrations of atoms at sensors. In this noise model, Eq. 1 becomes:
@@
where $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\sigma^{2}\boldsymbol{I})$, and $\sigma$ is the standard deviation. The other noise model is photon noise, which arises from the statistical nature of photons [35]. In most prior research, this is approximated as Poisson noise, and the measurement equations associated with this noise model typically follow a Poisson distribution:
@@
where $y_{i}=\sum_{j=1}^{N}M_{ij}x_{j}$ [6]. It should be noted that the constraint 2 in Eq. 3 prohibits negative entries in $\boldsymbol{M}$. While Poisson noise is the best way to approximate photon noise at this stage, its limitations for this project extend beyond just physical inaccuracies. The nonlinearity of photon noise adds complexities to the design and optimization of affected systems, and the use of Poisson random functions eliminates gradients, making it difficult to optimize the hardware components in the end-to-end model. To address these shortcomings, this project focuses on the signal dependency of photon noise and adopts a quasi-classical model known as MLGAUSS that uses a Gaussian variable with mean-variance equivalence [36], formulated as
[…]
It substitutes conventional Poisson approximations whenever a gradient computation is required, yet the term ”Poisson” is retained for consistency with other classical models involving no hardware optimization, facilitating straightforward comparisons. While this approach represents an alternative version of approximated photon noise besides Poisson noise, it is utilized exclusively during model training. For model testing, all models employ the Poisson noise model, as gradient computation is no longer necessary.
@@ IV-A Model Configuration
All coding strategies in this project for classification tasks share the same fundamental model structure, as illustrated in Fig. 2. In this project, we apply our model to handwritten number classification with the MNIST dataset and classification of hyperspectral data. To classify handwritten numbers using optical devices with limited photon budget, we built a PyTorch-based pipeline consisting of a scanner module including $\boldsymbol{M}$ and a noise generator to simulate the acquisition of $\tilde{\boldsymbol{y}}$ described in Eq. 2 and Eq. 3, and a classifier module classifying the measured objects by $\tilde{\boldsymbol{y}}$. For the classification task, we utilized a two-hidden layer artificial neural network (ANN) with 40 and 128 nodes, serving as the primary model. To prevent overfitting and enhance generalizability, we applied dropout after each hidden layer, followed by a rectified linear unit (ReLU) activation layer. Once operational, this model can serve various purposes, such as object classification and signal recovery. For instance, it can process images of handwritten numbers, providing the likelihood of each image corresponding to a specific number by cross-entropy loss. Furthermore, the classifier can function as a decoder to recover the input image by MSE loss, which is adopted in section IV-B.
During model training, the scanner computes the photon counts with noise and then transmits these data to the classifier. The noise is sampled from a predefined noise distribution and varies across epochs. Once the loss has been computed, the gradient descent is used to optimize the classifier parameters. In addition, we propose an end-to-end optimization model called Optical Neural Networks (ONN). What sets the ONN apart from other coding schemes is the learnable $\boldsymbol{M}$ which is not predefined or fixed. In other words, backpropagation is extended to the scanner, allowing the simultaneous optimization of the masks and the classifier under different noise models. This approach offers distinct advantages compared to traditional methods, which typically optimize the classifier independently of the scanner or optimize the scanner without the appropriate noise model.
@@
Optimizing the scanner can be challenging, as the Poisson random function typically lacks a gradient. We employ the reparameterized-Gaussian model mentioned in section III-D to address this limitation. This approach enables us to employ the PyTorch built-in reparameterization method, that is, rsample, and optimize the scanner more effectively under Poisson noise. During the training stages, variance plays a crucial role in the gradient computation, marking a significant difference compared to training under conventional Poisson models.